A 90.5kg man standing on a surface of negligible friction kicks forward a 45gm stone lying at his feet so that it acquire a speed of 3.05m/s. What velocity does the man acquire as a result?
1
Expert's answer
2021-08-16T08:35:56-0400

Given:

"M=90.5\:\rm kg"

"m=0.045\:\rm kg"

"u=3.05\:\rm m/s"


The law of conservation of momentum says

"0=Mv+mu"

Hence, the velocity of the man

"v=-\frac{mu}{M}=-\frac{0.045*3.05}{90.5}=-0.0015\:\rm m/s"